OpenStudy (vanz):

how do you solve this equation? 16x^2=64 ?

OpenStudy (anonymous):

2^4 x^2 = 2^6 x^2 = 2^2 x = 2 x = -2

OpenStudy (anonymous):

@saifoo it should be \[x = \pm \sqrt{4}\]

OpenStudy (saifoo.khan):

Yes, it will be -2 and 2.

OpenStudy (vanz):

Thank you guys! if you could look at this question I would highly appreciate, too. :D 6x^2\[6x ^{2}+11x-2=0\]

OpenStudy (saifoo.khan):

x = -2 x = 1/6

OpenStudy (vanz):

would you mind showing me the work if its not too much? :)

OpenStudy (saifoo.khan):

6x^2 + 11x -2 = 0 6x^2 +12x - x - 2 = 0 6x(x+2) -1(x+2) = 0 (x+2)(6x-1)=0 x= -2 x = 1/6